OpenStudy (anonymous):

Evaluate by substituting and simplifying |4.2 − 3x| − 6.1, if x = 2. Answer (A) -7.9 (B) -4.3 (C) 0.1 (D) -6.9

OpenStudy (anonymous):

| | of a negative number is positive ex: |-6| = 6

OpenStudy (anonymous):

-4.3

OpenStudy (anonymous):

u can see in my link

OpenStudy (anonymous):

yeah it will be |4.2-6|-6.1 |-1.8|-6.1 =1.8-6.1 =-4.3

OpenStudy (anonymous):

Thanks you guys ..your the best